OpenStudy (anonymous):

Can anybody please help me wit this? Simplify. The result should involve exactly one ratio.

OpenStudy (anonymous):

(1/x - x/x-1) /(2/x+1 - x-1/x)

OpenStudy (loser66):

\[\huge\dfrac{{\frac{1}{x}-\frac{x}{x-1}}}{{\frac{2}{x+1}-\frac{x-1}{x}}}\]

OpenStudy (anonymous):

can you help simplify it?

OpenStudy (loser66):

just make it clear from your form. don't know how to do yet. XD. Do you have any idea?

OpenStudy (loser66):

No? so, we can do together, separate and calculate the numerator first, make them have the same denominator, what do you get?

OpenStudy (anonymous):

not really :/

OpenStudy (anonymous):

I would clear the denominators what is the LCM of the two denominators?

OpenStudy (anonymous):

x+1?

OpenStudy (loser66):

@RH, show me your work. I don't know how to do for the whole thing, just step by step. We can carefully check each of them, OK?

OpenStudy (anonymous):

x+1/ x(x+1) - xˆ2+1 / x(x+1)

OpenStudy (anonymous):

is this correct? I am not sure :(

OpenStudy (anonymous):

Looking good so far.

OpenStudy (anonymous):

So RH what do you get when you combined the two fractions in the numerator?

OpenStudy (anonymous):

i dont know:(

OpenStudy (anonymous):

Well let's focus on the numerator ok...

OpenStudy (anonymous):

So that is\[\frac{ 1 }{ x }-\frac{ x }{ x-1 }\] What is the LCD of those two fractions?

OpenStudy (anonymous):

x+1?

OpenStudy (anonymous):

No think about normal fractions what is the LCD of \[\frac{ 2 }{ 3 }-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Right ... now how did you get that?

OpenStudy (anonymous):

so you would have to multiply 2/3 by 2/2 and 1/2 by 3/3 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Paynesdad

OpenStudy (anonymous):

So now to this problem what would the LCD of the top problem be using the same thought process

OpenStudy (anonymous):

x(x+1) ?

OpenStudy (anonymous):

Close what are the two denominators in the top part of the ratio?

OpenStudy (anonymous):

x and x+1

OpenStudy (anonymous):

x+1 is a denominator in the top part of the ratio? \[\frac{ 1 }{ x }-\frac{ x }{ x-1}\]

OpenStudy (anonymous):

1 and x, sorry

OpenStudy (anonymous):

No those are the numerators, what are the denominators?

OpenStudy (anonymous):

x and x-1

OpenStudy (anonymous):

There you go...so what does the top and bottom of the first fraction need to be multiplied by?